# Supervised-Contrastive Loss Learns Orthogonal Frames and Batching Matters

### v1

@@ Abstract
Supervised contrastive loss (SCL) is a competitive and often superior alternative to the cross-entropy (CE) loss for classification. In this paper we ask: what differences in the learning process occur when the two different loss functions are being optimized? To answer this question, our main finding is that the geometry of embeddings learned by SCL forms an orthogonal frame (OF) regardless of the number of training examples per class. This is in contrast to the CE loss, for which previous work has shown that it learns embeddings geometries that are highly dependent on the class sizes. We arrive at our finding theoretically, by proving that the global minimizers of an unconstrained features model with SCL loss and entry-wise non-negativity constraints form an OF. We then validate the model’s prediction by conducting experiments with standard deep-learning models on benchmark vision datasets. Finally, our analysis and experiments reveal that the batching scheme chosen during SCL training plays a critical role in determining the quality of convergence to the OF geometry. This finding motivates a simple algorithm wherein the addition of a few binding examples in each batch significantly speeds up the occurrence of the OF geometry.
Code available here.
@@ 1 Introduction
The prevalence of deep-neural networks (DNNs) has led to a growing research interest in understanding their underlying mechanisms. A recent research thread, focusing on classification tasks, explores whether it is possible to describe the structure of weights learned by DNNs when trained beyond zero-training error. The specific characteristics of this structure will depend on the DNN being used, the dataset being trained on, and the chosen optimization hyperparameters. However, can we identify macroscopic structural characteristics that are common among these possibilities?
In an inspiring study, [PHD20] demonstrates this is possible for the classifiers and for the embeddings when training with cross-entropy (CE) loss. Through extensive experiments over multiple architectures and balanced datasets with an equal number of examples per class, they found that the geometries of classifiers and of centered class-mean embeddings (i.e., outputs of the last hidden layer) consistently converge during training to a common simplex equiangular tight frame (ETF), a structure composed of vectors that have equal norms and equal angles between them, with the angles being the maximum possible. Moreover, neural-collapse (NC), a property where embeddings of individual examples from each class converge to their class-mean embedding, was also observed. Numerous follow-up studies have delved deeper into explaining this phenomenon and further investigating how the converging geometry changes with class imbalances. The unconstrained-features model (UFM), proposed by [MPP20, Fan+21, Gra+21, Ji+21, LS20, TB22], plays a central role in the majority of these follow-up studies. Specifically, the UFM serves as a theoretical abstraction for DNN training, in which the network architecture is viewed as a powerful black-box that generates embeddings without any restrictions in the last (hidden) layer. For CE loss, the UFM minimizes $\min_{\mathbf{w}_{c}\in\mathbb{R}^{d},\mathbf{h}_{i}\in\mathbb{R}^{d}}\mathcal{L}_{\text{CE}}\big{(}\{\mathbf{w}_{c}\}_{c\in[k]},\{\mathbf{h}_{i}\}_{i\in[n]}\big{)}$ in which both classifiers $\mathbf{w}_{c}$ for the $k$ classes and embeddings $\mathbf{h}_{i}$ for the $n$ training examples are unconstrainedly optimized over $\mathbb{R}^{d}$. [Zhu+21, Gra+21, Fan+21] have verified that the global minimum of this non-convex problem satisfies NC and follows the ETF geometry observed in DNN experiments by [PHD20]. Interestingly, recent works by [Thr+22, Fan+21, Beh+23] have demonstrated that the global optimum of the UFM changes when classes are imbalanced. However, the new solution still predicts the geometry observed in DNN experiments, providing evidence that, despite its oversimplification, the UFM is valuable in predicting structural behaviors.
@@
[…]
Figure 1: Experimental validation of our key finding: SCL learns orthogonal class-mean embeddings irrespective of class imbalances. Shown are box plots of pairwise angular distance $\Delta_{\text{ang}}=1-\nicefrac{{1}}{{\pi}}\arccos\big{(}\nicefrac{{\bm{\mu}_{c}^{\top}\bm{\mu}_{c^{\prime}}}}{{\|\bm{\mu}_{c^{\prime}}\|\|\bm{\mu}_{c}\|}}\big{)}$ between class-mean embeddings $\bm{\mu}_{c},\bm{\mu}_{c^{\prime}},c\neq c^{\prime}\in[k]$ of a ResNet-18 (bottom) and DenseNet-40 (top) trained with SCL on $R$-STEP-imbalanced CIFAR-10. In all three cases, our prediction highlighted in green and corresponding to an orthogonal angle is remarkably close to the measurements. Compare to the red line predicted by [Gra+21] for balanced data. No explicit characterization has been known previously for imbalanced data.
Expanding beyond the scope of CE minimization, [Gra+21] also used the UFM to determine whether optimizing with the supervised-contrastive loss (SCL) results in any alterations to the geometric structure of the learned embeddings. [^1] SCL is designed to train only embeddings and is an extension of the well-known contrastive loss used for unsupervised learning, adapted to supervised datasets [Kho+20]; see Equation (4). They proved for balanced datasets that the global solution of $\min_{\mathbf{h}_{i}\in\mathbb{R}^{d}}\mathcal{L}_{\text{SCL}}\big{(}\{\mathbf{h}_{i}\}_{i\in[n]}\big{)}$ remains a simplex ETF. This suggested that CE and SCL find the same embedding geometries. In order to verify this empirically, they trained a ResNet-18 on CIFAR10 and calculated the angular distance between the class-mean embeddings $\bm{\mu}_{c},\bm{\mu}_{c^{\prime}}$ for $c\neq c^{\prime}$. In Fig. 1(a), we have reproduced their experiment and presented a boxplot illustrating the distribution of the measured angular distances. Note that the angular distance value, as predicted by the analysis carried out in [Gra+21] and highlighted in red, shows a slight deviation from the measured values. Moreover, their prediction only applies to balanced data and no prior work has explicitly characterized the geometry of SCL loss with class imbalances. These observations give rise to the following research questions motivating our work:
When classes are balanced, what could account for the difference between the UFM prediction from [Gra+21] and the experimental values in Fig. 1(a)? Is the UFM unable to accurately predict the geometry learned by SCL? Additionally, what is the geometry of the embeddings under imbalances? Ultimately, is the implicit geometry of embeddings the same when optimizing with either the CE or the SCL loss?
1.1 Summary of contributions
This paper answers all the above questions.
Key finding. We find that SCL learns class-mean embeddings that are pairwise orthogonal, hence they form an orthogonal frame (OF), a structure composed of vectors that have equal norms and are mutually orthogonal to each other. Interestingly, this geometry remains consistent regardless of the label distribution, i.e., the number of examples per class. This finding highlights a notable distinction between SCL and CE loss, as the geometry of embeddings learnt with CE is known to be highly-influenced by the label distribution.
Experimental justification. An initial experimental validation of our finding is illustrated in Fig. 1. The experiments correspond to STEP-imbalanced CIFAR-10 data with five majority/minority classes and imbalance ratio $R=1,10,100$. The green line highlights the angular distance of orthogonal angles (corresponding to an OF), which is $1/2$, and remarkably aligns with the measured angles for all values of the imbalance ratio $R$. Extensive additional experiments for Long-tailed (LT) distributions, other datasets, and architectures, are presented in Sec. 3 and Sec. D in the appendix.
@@ 1.1 Summary of contributions
Theoretical justification. Our finding emerged through theoretical analysis by investigating the global minimizers of an extended version, denoted as $\text{UFM}_{+}$, of the original unconstrained-features model (UFM). This refinement accounts for the presence of ReLU activations in the last hidden layer of commonly used architectures by minimizing the SCL over the non-negative orthant. Concretely, we show that all global solutions of $\min_{\mathbf{h}_{i}\geq 0}\mathcal{L}_{\text{SCL}}\big{(}\{\mathbf{h}_{i}\}_{i\in[n]}\big{)}$ satisfy NC and the corresponding class-mean features form an OF. This explains the implicit geometry observed in our experiments, eliminating the requirement for a heuristic centering of the class-mean embeddings as implemented in the majority of prior studies.
@@
Analysis of batching and implications. Finally, we examine the impact of batching on determining the learned embedding geometry. By utilizing the $\text{UFM}_{+}$, we establish straightforward criteria for the batching scheme, ensuring that any global minimizer forms an OF. By utilizing these conditions, we provide an explanation for the observed substantial degradation in convergence towards an OF when employing a fixed batch partition instead of randomly reshuffling the batches at each training epoch. Intriguingly, we demonstrate that any arbitrary batching partition can be transformed into one that fulfills our theoretical conditions by introducing $k$ so-called “binding examples” in each batch. Moreover, as verified in Sec. 5.2, this approach consistently accelerates the convergence of the learned geometry towards an OF.
1.2 Organization
The remaining portion of the paper is organized as follows. In Sec. 2, we formally introduce the SCL and provide an overview of the OF geometry. Sec. 3 presents and discusses the experimental evidence that supports our main finding. The theoretical justification for our finding is presented in Sec. 4, where we analyze both full-batch and mini-batch SCL scenarios. Building upon the insights gained from the analysis of mini-batch SCL, Sec. 5 explores the critical role of batching and proposes a binding-examples scheme aimed at enhancing the convergence quality of the implicit geometry. To position our results in the context of related works, Sec. 6 further discusses the most closely related studies and presents a detailed comparison between the $\text{UFM}_{+}$ model, which we rely on in this paper, and the previously-used UFM. Sec. 7 concludes the paper by reflecting on the contribution of our findings and discussing their implications at the intersection of the literature on neural-collapse phenomena and contrastive-type losses. The appendix contains all the proofs, along with additional experimental details and supplementary experiments.
Notation.  For a positive integer $n$, we denote $[n]:=\{1,2,3,...,n\}$. For matrix $\mathbf{V}\in\mathbb{R}^{m\times n}$, $\mathbf{V}[i,j]$ denotes its $(i,j)$-th entry, $\mathbf{v}_{j}$ denotes the $j$-th column, $\mathbf{V}^{\top}$ its transpose. We denote $\|\mathbf{V}\|_{F}$ the Frobenius norm of $\mathbf{V}$. We use $\mathbf{V}\propto\mathbf{X}$ whenever the two matrices are equal up to a scalar constant. We use $\mathbf{V}\geq 0$ to denote the entry-wise non-negativity of the elements of $\mathbf{V}$, i.e., $\mathbf{V}[i,j]\geq 0,\forall i\in[m],j\in[n]$. Finally, $\mathds{I}_{m}$ denotes the identity matrix of size $m$.
@@ 2 Setup
Consider $k$-class classification and training set $S=\{(\mathbf{x}_{i},y_{i}):i\in[n]\}$ where $\mathbf{x}_{i}\in\mathbb{R}^{d},\,y_{i}\in[k]$ represent data samples and corresponding class labels. We use $n_{c}$ to denote the number of examples in class $c\in[k]$, and $\mathbf{h}_{\bm{\theta}}(\cdot)$ to denote the last-layer feature-embeddings of a deep-net parameterized by ${\bm{\theta}}$. We compute the SCL on a given batch $B\subseteq[n]$ of training examples as follows,
@@
where $\tau$ is a positive scalar temperature hyper-parameter, and $n_{B,y_{i}}$ is the number of examples in batch $B$ belonging to class $c=y_{i}$. To train a deep-net, we minimize SCL (3) over the network parameters ${\bm{\theta}}$ on a set of batches $\mathcal{B}$ chosen from the training set. As introduced by [Kho+20, ], SCL requires normalized features, so we assume $\lVert\mathbf{h}_{{\bm{\theta}}}(\mathbf{x}_{i})\rVert=1$. Furthermore, throughout this paper, we assume that $d\geq k$ and $n_{c},n_{B,c}\geq 2,c\in[k]$. [^2] When using SCL, it is common practice to add an augmented version of the datapoints in each batch to itself [Kho+20, Gra+21]. This practice, referred to as batch duplication [Gra+21], helps ensure that each datapoint in a batch has at least one other example in the same class to compare to during training.
[…]
Figure 2: ResNet-18 model trained on imbalanced datasets with SCL. Convergence of last-layer embeddings to the OF geometry as measured by $\Delta_{\mathbf{G}_{\mathbf{M}}}:=\|\frac{\mathbf{G}_{\mathbf{M}}}{\|\mathbf{G}_{\mathbf{M}}\|_{F}}-\frac{\mathds{I}_{k}}{\|\mathds{I}_{k}\|_{F}}\|_{F}$, where $\mathbf{G}_{\mathbf{M}}=\mathbf{M}^{\top}\mathbf{M}$ is the Gram matrix of class-mean embeddings. $\Delta_{\mathbf{G}_{\mathbf{M}}}$ is consistently on the order of $0.01$ regardless of dataset or imbalance.
We let $\textbf{H}_{d\times n}=\begin{bmatrix}\textbf{h}_{1},\cdots,\textbf{h}_{n}\end{bmatrix}$ denote an embeddings matrix where each column corresponds to one of the examples in the training set, i.e., $\mathbf{h}_{i}:=\mathbf{h}_{\bm{\theta}}(\mathbf{x}_{i})$. The embeddings geometry or so-called implicit geometry 333 The specific terminology is adopted from [Beh+23]. refers to the norms and pairwise-angles of these vectors $\mathbf{h}_{i},i\in[n]$. Note these quantities correspond exactly to the entries of the Gram matrix ${\mathbf{H}}^{\top}{\mathbf{H}}$. In order to characterize the geometry, we need the following definitions.
@@
When the within-class variation of embeddings is negligible, i.e., ${\mathbf{H}}$ satisfies NC, it suffices to focus on the class-mean embeddings ${\bm{\mu}}_{c}=\frac{1}{n_{c}}\sum_{i:y_{i}=c}\mathbf{h}_{i},c\in[k]$ and their respective matrix $\textbf{M}_{d\times k}=\begin{bmatrix}{\bm{\mu}}_{1},\cdots,{\bm{\mu}}_{k}\end{bmatrix}$. That is, it suffices to study $\mathbf{G}_{\mathbf{M}}=\mathbf{M}^{\top}\mathbf{M}$ instead of ${\mathbf{H}}^{\top}{\mathbf{H}}.$ With these, we can formally define the OF geometry for embeddings ${\mathbf{H}}$ as follows.
@@
3 SCL learns OF geometries: Empirical findings

### v2

# Symmetric Neural-Collapse Representations with Supervised Contrastive Loss: The Impact of ReLU and Batching

@@ Abstract
Supervised contrastive loss (SCL) is a competitive and often superior alternative to the cross-entropy loss for classification. While prior studies have demonstrated that both losses yield symmetric training representations under balanced data, this symmetry breaks under class imbalances. This paper presents an intriguing discovery: the introduction of a ReLU activation at the final layer effectively restores the symmetry in SCL-learned representations. We arrive at this finding analytically, by establishing that the global minimizers of an unconstrained features model with SCL loss and entry-wise non-negativity constraints form an orthogonal frame. Extensive experiments conducted across various datasets, architectures, and imbalance scenarios corroborate our finding. Importantly, our experiments reveal that the inclusion of the ReLU activation restores symmetry without compromising test accuracy. This constitutes the first geometry characterization of SCL under imbalances. Additionally, our analysis and experiments underscore the pivotal role of batch selection strategies in representation geometry. By proving necessary and sufficient conditions for mini-batch choices that ensure invariant symmetric representations, we introduce batch-binding as an efficient strategy that guarantees these conditions hold.
@@ 1 Introduction
The prevalence of deep-neural networks (DNNs) has led to a growing research interest in understanding their underlying mechanisms. A recent research thread, focusing on classification tasks, explores whether it is possible to describe the structure of weights learned by DNNs when trained beyond zero-training error. The specific characteristics of this structure will depend on the DNN being used, the dataset being trained on, and the chosen optimization hyperparameters. Yet, is it possible to identify macroscopic structural characteristics that are common among these possibilities?
In an inspiring study, [PHD20] demonstrates this is possible for the classifiers and for the embeddings when training with cross-entropy (CE) loss and balanced datasets. Through extensive experiments over multiple architectures and datasets with an equal number of examples per class, they found that the geometries of classifiers and of centered class-mean embeddings (outputs of the last hidden layer) consistently converge during training to a common simplex equiangular tight frame (ETF), a structure composed of vectors that have equal norms and equal angles between them, with the angles being the maximum possible. Moreover, they observed neural-collapse (NC), a property where embeddings of individual examples from each class converge to their class-mean embedding.
Numerous follow-up studies have delved deeper into explaining this phenomenon and further investigating how the converging geometry changes with class imbalances. The unconstrained-features model (UFM), proposed independently by [MPP20, Fan+21, Gra+21, LS20], plays a central role in the majority of these follow-up studies. Specifically, the UFM serves as a theoretical abstraction for DNN training, in which the network architecture is viewed as a powerful black-box that generates embeddings without any restrictions in the last (hidden) layer. For CE loss, the UFM minimizes $\min_{\mathbf{w}_{c}\in\mathbb{R}^{d},\mathbf{h}_{i}\in\mathbb{R}^{d}}\mathcal{L}_{\text{CE}}\big{(}\{\mathbf{w}_{c}\}_{c\in[k]},\{\mathbf{h}_{i}\}_{i\in[n]}\big{)}$ in which both classifiers $\mathbf{w}_{c}$ for the $k$ classes and embeddings $\mathbf{h}_{i}$ for the $n$ training examples are unconstrainedly optimized over $\mathbb{R}^{d}$. [Zhu+21, Gra+21, Fan+21] have verified that the global minimum of this non-convex problem satisfies NC and follows the ETF geometry observed in DNN experiments by [PHD20]. Recent works by [Thr+22, Fan+21, Beh+23] have demonstrated that the global optimum of the UFM changes when classes are imbalanced. Yet, the new solution still predicts the geometry observed in DNN experiments, providing evidence that, despite its oversimplification, the UFM is valuable in predicting structural behaviors.
Expanding beyond the scope of CE minimization, [Gra+21] also used the UFM to determine whether optimizing with the supervised-contrastive loss (SCL) results in any alterations to the geometric structure of the learned embeddings. [^1] SCL is designed to train only embeddings and is an extension of the well-known contrastive loss used for unsupervised learning, adapted to supervised datasets [Kho+20]; see Equation (4).
@@
Figure 1: Gram matrices, $\mathbf{G}_{\mathbf{M}}/\max_{i,j}|\mathbf{G}_{\mathbf{M}}[i,j]|$, of class-means of learned feature embeddings at last epoch of training, with ResNet-18 on MNIST. SCL+ReLU: the mean feature embeddings for different classes are mutually orthogonal, forming an OF, regardless of imbalance (imbalance ratio $R=1,10,100$). This invariance does not hold in the absence of ReLU for SCL. Further, CE feature geometry is also sensitive to imbalance. The label distribution is STEP imbalanced, with the first five classes as majorities and the rest as minorities. See Sec.E.2 in the appendix for more information.
They proved for balanced datasets that the global solution of $\min_{\mathbf{h}_{i}\in\mathbb{R}^{d}}\mathcal{L}_{\text{SCL}}\big{(}\{\mathbf{h}_{i}\}_{i\in[n]}\big{)}$ remains a simplex ETF, suggesting that CE and SCL find the same embedding geometries. In Fig. 1, we investigate the geometry of class-means in presence of class-imbalance. The prediction in [Gra+21] only applies to balanced data (SCL with $R=1$ in Fig. 1) and no prior work has explicitly characterized the geometry of SCL with class imbalances. Observing the middle column of the figure, note that the geometry of embeddings changes drastically [^2] Analogous finding is reported by [Zhu+22], although not visualized as done here. as the imbalance ratio $R$ increases. This behavior of SCL is consistent with the CE embeddings geometry, which, as discussed previously, also changes with the imbalance; see last column of Fig. 1.
In this paper, we arrive at a surprising finding: simply introducing a ReLU activation at the final layer restores the symmetry in SCL-trained representations. This phenomenon is clearly illustrated in the first column of Fig. 1. Below is a detailed description of our contributions.
1.1 Summary of contributions
We conduct an in-depth examination of the geometry of training representations (embeddings) of SCL, particularly in relation to varying levels of dataset imbalances. Our study leads us to identify and capture, both analytically and empirically, and for the first time, the impact on the representation geometry of: (i) a straightforward architectural modification, specifically the incorporation of a ReLU activation at the final layer, and (ii) the batch-selection strategy.
Impact of ReLU. We find that the addition of a ReLU activation to the last-layer of the architecture results in an orthogonal frame (OF) geometry. That is, a geometry structure composed of class-mean embedding vectors that have equal norms and are mutually orthogonal to each other. A preliminary experimental validation of this finding is illustrated in Fig. 1. The experiments correspond to STEP-imbalanced MNIST data with five majority/minority classes and imbalance ratio $R=1,10,100$. Each heatmap represents the pairwise inner products between the class-means of learned feature embeddings. The figure (first column) reveals that with the addition of ReLU at the last layer, SCL learns orthogonal features regardless of class imbalance. This is in stark contrast to the varying geometries of vanilla SCL (optimized commonly without ReLU) and CE loss. Extensive additional experiments for Long-tailed (LT) distributions, other datasets and architectures, are presented in Sec. 3 and also Sec. E in the appendix. We also present experimental findings that confirm ReLU symmetrises the geometry without compromising test accuracy.
@@ 1.1 Summary of contributions
Figure 2: Distance of learned embeddings to the OF geometry as measured by $\|\frac{\mathbf{G}_{\mathbf{M}}}{\|\mathbf{G}_{\mathbf{M}}\|_{F}}-\frac{\mathds{I}_{k}}{\|\mathds{I}_{k}\|_{F}}\|_{F}$ where $\mathbf{G}_{\mathbf{M}}=\mathbf{M}^{\top}\mathbf{M}$ is the Gram matrix of class-mean embeddings. For MNIST and CIFAR10 we use ResNet-18 and for the more complex CIFAR100 and Tiny-ImageNet, we use ResNet-34 and train using batch-binding (see Sec. 5). Regardless of the imbalance level $R$, the embeddings consistently converge to the OF geometry.
@@
We support this finding by theoretically investigating the global minimizers of an extended version, denoted as $\text{UFM}_{+}$, of the original unconstrained-features model (UFM). This refinement accounts for the presence of ReLU activations which we impose on the last-layer by minimizing SCL over the non-negative orthant. Concretely, we show that all global solutions of $\min_{\mathbf{h}_{i}\geq 0}\mathcal{L}_{\text{SCL}}\big{(}\{\mathbf{h}_{i}\}_{i\in[n]}\big{)}$ satisfy NC and the corresponding class-mean features form an OF. This explains the convergence of feature geometry to an OF as consistently observed in our experiments; for example, see Fig. 2.
@@
Batching and its implications. Furthermore, we identify the crucial impact of batch selection strategies during SCL training in shaping the learned embedding geometry. Concretely, by analyzing the $\text{UFM}_{+}$, we establish straightforward criteria for the batching scheme, ensuring that any global minimizer of $\text{UFM}_{+}$ forms an OF. These conditions explain the substantial degradation in convergence towards an OF when employing a fixed batch partition instead of randomly reshuffling the batches at each training epoch. Intriguingly, we demonstrate that any arbitrary batching partition can be transformed into one that fulfills our theoretical conditions by performing so-called batch-binding. Through extensive experiments in Sec. 5.2 and in the appendix, we show that batch-binding consistently accelerates the convergence of the learned geometry towards an OF and even improves convergence to ETF geometry under balanced training in the absence of ReLU.
@@ 2 Setup
Notation.  For a positive integer $n$, $[n]:=\{1,2,3,...,n\}$. For matrix $\mathbf{V}\in\mathbb{R}^{m\times n}$, $\mathbf{V}[i,j]$ denotes its $(i,j)$-th entry, $\mathbf{v}_{j}$ denotes the $j$-th column, $\mathbf{V}^{\top}$ its transpose. We denote $\|\mathbf{V}\|_{F}$ the Frobenius norm of $\mathbf{V}$. We use $\mathbf{V}\propto\mathbf{X}$ whenever the two matrices are equal up to a scalar constant. We use $\mathbf{V}\geq 0$ to denote the entry-wise non-negativity, i.e., $\mathbf{V}[i,j]\geq 0,\forall i\in[m],j\in[n]$. Finally, $\mathds{I}_{m}$ denotes the identity matrix of size $m$.
Consider $k$-class classification and training set $S=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{n}$ where $\mathbf{x}_{i}\in\mathbb{R}^{d},\,y_{i}\in[k]$ represent data samples and corresponding class labels. We use $n_{c}$ to denote the number of examples in class $c\in[k]$, and $\mathbf{h}_{\bm{\theta}}(\cdot)$ to denote the last-layer feature-embeddings of a deep-net parameterized by ${\bm{\theta}}$. We compute the SCL on a given batch $B\subseteq[n]$ of training examples as follows,
@@
where $\tau$ is a positive scalar temperature hyper-parameter, and $n_{B,y_{i}}$ is the number of examples in batch $B$ belonging to class $c=y_{i}$. To train a deep-net, we minimize SCL (3) over the network parameters ${\bm{\theta}}$ on a set of batches $\mathcal{B}$ chosen from the training set. As introduced by [Kho+20, ], SCL requires normalized features, so we assume $\lVert\mathbf{h}_{{\bm{\theta}}}(\mathbf{x}_{i})\rVert=1$. Furthermore, we assume that $d\geq k$ and $n_{c},n_{B,c}\geq 2,c\in[k]$. [^3] When using SCL, it is common practice to add an augmented version of the datapoints in each batch to itself [Kho+20, Gra+21]. This practice, referred to as batch duplication, helps ensure that each datapoint in a batch has at least one other example in the same class to compare to during training. We let $\textbf{H}_{d\times n}=\begin{bmatrix}\textbf{h}_{1},\cdots,\textbf{h}_{n}\end{bmatrix}$ denote an embeddings matrix where each column corresponds to one of the examples in the training set, i.e., $\mathbf{h}_{i}:=\mathbf{h}_{\bm{\theta}}(\mathbf{x}_{i})$. The embeddings geometry or so-called implicit geometry444 The specific terminology is adopted from [Beh+23]. refers to the norms and pairwise-angles of these vectors $\mathbf{h}_{i},i\in[n]$. Note these quantities correspond exactly to the entries of the Gram matrix ${\mathbf{H}}^{\top}{\mathbf{H}}$. To characterize the geometry, we need the following definitions.
@@
When the within-class variation of embeddings is negligible, i.e., ${\mathbf{H}}$ satisfies NC, it suffices to focus on the class-mean embeddings ${\bm{\mu}}_{c}=\frac{1}{n_{c}}\sum_{i:y_{i}=c}\mathbf{h}_{i},c\in[k]$ and the respective matrices $\textbf{M}_{d\times k}=\begin{bmatrix}{\bm{\mu}}_{1},\cdots,{\bm{\mu}}_{k}\end{bmatrix}$ and $\mathbf{G}_{\mathbf{M}}=\mathbf{M}^{\top}\mathbf{M}$ instead of ${\mathbf{H}}$ and ${\mathbf{H}}^{\top}{\mathbf{H}}.$ With these, we can formally define the OF geometry for embeddings.
@@
3 SCL with ReLU learns OF geometries: Empirical findings
Experimental setup.  Following the setup of [Gra+21, ], our models consist of a backbone (ResNet, DenseNet, etc) with a normalizing layer to output features with unit norm. For all geometric convergence results, we apply a ReLU on output features before normalization. We employ Stochastic Gradient Descent (SGD) with a learning rate of $0.1$, momentum set to $0.9$ and with no weight decay. Furthermore, as in [Gra+21, Kho+20, ], we set the temperature parameter $\tau=0.1$ as [Kho+20, ] have found that it yields optimal performance. In addition, we empirically find that convergence to OF is not highly dependent on $\tau$ for values near $0.1$, yet the specific choice affects the speed of convergence. Code is available here.